### v1

# Global Prompt Refinement with Non-Interfering Attention Masking for One-Shot Federated Learning

Abstract
Federated Prompt Learning (FPL) enables communication-efficient adaptation by tuning lightweight prompts on top of frozen pre-trained models. Existing FPL methods typically rely on global information, which is only available after the second training round, to facilitate collaboration among client models. Therefore, they are inherently dependent on multi-round communication to fully exhibit their strengths. Moreover, existing one-shot federated learning methods typically focus on fitting seen tasks, but lack cross-task generalization. To bridge this gap, we propose the Global Prompt Refinement with Non-Interfering Attention Masking (GPR-NIAM) method for one-shot FPL. The core idea is to design a masking mechanism that restricts excessive interaction between the original text embeddings and the learnable prompt embeddings. GPR-NIAM achieves this through the collaboration of two key modules. Firstly, the attention isolation module suppresses attention from the learnable prompt tokens to the original text tokens, and reweights the reverse attention which preserves generalization across tasks. Secondly, the cross-silo collaborative refinement module integrates decentralized visual knowledge into a unified base and calibrates the global prompt through multi-source cross-modal knowledge alignment, further mitigating the inconsistency caused by data heterogeneity. Extensive experiments conducted on ten benchmark datasets under two tasks show that GPR-NIAM outperforms eight state-of-the-art methods in both class-level and domain-level generalization.
1 Introduction
Federated learning (FL) enables collaborative modeling across data sources in a privacy-preserving manner. It learns global knowledge by aggregating model weights or gradients instead of moving data to a central entity mcmahan2017communication ; cai2024lgfgad ; 9820618 ; yi2024federated ; yi2023pfedes ; 11091509 . Recently, Federated Prompt Learning (FPL) has emerged to support parameter-efficient fine-tuning in this setting. It not only leverages pre-trained models (e.g., CLIP radford2021learning ) more effectively, but also reduces communication overhead between FL clients and the FL server by transmitting only learnable prompts qi2025federated ; zhao2023fedprompt ; ren2025advances ; fan2025ten . However, heterogeneous data distributions across clients often degrade model aggregation performance and slow down convergence yang2023efficient ; bai2024diprompt ; li2020federated ; guo2025federated ; guo2024sample . This is because significant differences in the update directions across local models can cause conflicts during aggregation pan2024federated ; cui2024harmonizing ; huang2024federated ; dong2022federated_FCIL ; fu2025federated . As a result, existing FPL methods still require multi-round communications between clients and the server to reconcile inconsistencies arising from client-specific training.
To deal with data heterogeneity, early FPL solutions (e.g., PromptFoilo pan2024federated and DP-FPL tran2025privacy ) typically focused on incorporating global prompt information to constrain the update direction of local prompts in order to enhance the consistency and stability of federated optimization pan2024federated ; cui2024harmonizing ; tran2025privacy ; li2024global . However, since global prompt information is typically available only after the second communication round, these methods inherently rely on multi-round communications. This poses a challenge in real-world scenarios with limited communication resources. To this end, one-shot federated learning (OSFL), which requires only a single round of client-server interaction, offers a promising direction for further exploration zhang2022dense ; zeng2024one ; diao2023towards ; yang2024exploring ; cai2024fgad . OSFL significantly reduces the communication overhead by eliminating iterative exchanges, making it attractive for real-world deployments with strict latency or bandwidth constraints yao2025fedmho ; ma2025geometric ; hu2024fedmut . However, directly applying existing OSFL methods to FPL settings is not suitable, as they are typically optimized for seen tasks while neglecting the preservation of generalization to unseen tasks.
Figure 1: An overview of the proposed GPR-NIAM method. It supports three interaction types: full attention (0), hard masking (-$\infty$), and reweighting ($\lambda$). Hard masking blocks learnable tokens $T_{P}$ from affecting original text tokens $T_{T}$, while $\lambda$-weighted links enable partial influence from special tokens $T_{EOS}$. The upper triangular mask follows causal attention for autoregressive decoding.
To bridge this important gap, we propose the Global Prompt Refinement with Non-Interfering Attention Masking (GPR-NIAM) method. It prevents the modification of the original token embeddings through a masking mechanism, thereby preserving transferable knowledge across tasks. As shown in Figure 1, the interaction between the learnable prompt embeddings and the original text token embeddings is unidirectional, which ensures that the latter remain unchanged. Specifically, GPR-NIAM consists of two main modules. Firstly, the Attention Isolation module selectively regulates the flow of prompt information through a novel masking strategy. It suppresses attention from the learnable prompt tokens to the original text tokens across all encoder layers, thereby effectively preventing alterations to the textual input token embeddings. In addition, it re-weights the reverse attention to encourage the model to extract task-relevant features more effectively from the original texts. Secondly, the cross-silo collaborative refinement (CSCR) module leverages visual representations extracted from multiple clients and utilizes multi-source visual supervision to refine the global prompt in a centralized manner. This further mitigates the negative impact of data heterogeneity.
We conduct extensive experiments on 10 datasets under two tasks: 1) base-to-base/novel generalization and 2) leave-one-domain-out generalization, and carry out performance comparisons, ablation studies, in-depth analysis and case studies. The results demonstrate that GPR-NIAM achieves superior overall performance by maintaining a strong balance between fitting seen tasks and generalizing to unseen ones, significantly outperforming eight competitive state-of-the-art approaches.
2 Related Work
2.1 Federated Prompt Learning
FPL aims to integrate FL with prompt learning by freezing the backbone model within the FL framework and introducing a set of learnable prompt parameters pan2024federated ; cui2024harmonizing ; tran2025privacy ; li2024global ; su2024federated ; jia2025llm ; guo2023pfedprompt ; deng2024unlocking . This enables efficient personalization for on-device data while significantly reducing communication overhead, as only prompt parameters are shared between clients and the server. To mitigate the challenges posed by data heterogeneity across clients, popular FPL methods (e.g., FedOTP li2024global , FedPGP cui2024harmonizing , PromptFolio pan2024federated , DP-FPL tran2025privacy ) mostly adopt a shared prompt paradigm. In this design, a global prompt is constructed on the server and combined with client-specific information to improve training stability and generalization under non-IID distributions. However, these methods are essentially designed for multi-round communications and require access to global feedback, which limits their effectiveness in one-shot FL scenarios.
2.2 One-Shot Federated Learning
OSFL is a communication-efficient variant of federated learning, where model training is completed within a single round of client-server interaction zhang2022dense ; zeng2024one ; li2025feature ; diao2023towards ; yang2024exploring ; yang2024feddeo ; allouah2024revisiting ; chiang2023optimal ; heinbaugh2023data ; wang2024one . Recent studies on OSFL mainly focus on generative and ensemble-based approaches. The former employs generative models to synthesize proxy data or representations on the server. For example, FedCVAE heinbaugh2023data uses a variational autoencoder to compress local knowledge. FedDISC yang2024exploring and FedDEO yang2024feddeo leverage diffusion models to generate synthetic training data. The latter is dedicated to refining the fusion of model predictions from different clients. For example, FedOV diao2023towards introduces an open-set recognition-based voting mechanism to address label skew in OSFL; IntactOFL zeng2024one learns weighted or expert-based combinations to enhance overall performance. However, these methods are only designed for conventional tasks, aiming to fit the training objective while overlooking the preservation of cross-task generalization.
3 Preliminaries
Suppose there are $K$ clients denoted by $\mathcal{C}=\{C_{1},C_{2},\dots,C_{K}\}$, and each client $C_{k}\in\mathcal{C}$ holds a private dataset $\mathcal{D}_{k}$. The goal of traditional FL is to collaboratively train a global model $f$ by minimizing the objective: $\min_{f}\sum_{k=1}^{K}p_{k}\cdot\mathcal{L}(f,\mathcal{D}_{k})$, where $p_{k}=\frac{|\mathcal{D}_{k}|}{\sum_{j=1}^{K}|\mathcal{D}_{j}|}$ is the weight of client $k$, and $\mathcal{L}(\cdot)$ is a task-specific loss function. In FPL, each client $k$ optimizes a set of learnable prompts $\delta_{k}$ using its private dataset $\mathcal{D}_{k}$, while keeping the backbone model $f$ frozen: $\min_{\delta_{k}}\mathcal{L}(f;\delta_{k},\mathcal{D}_{k})$. To support class-level prediction, the prompt for class $c$ is constructed by concatenating the learnable prompt and a fixed class token. Specifically, let $\mathbf{p}_{c}\in\mathbb{R}^{N_{t}\times d_{t}}$ denote the class token for class $c$, and $\delta_{k}\in\mathbb{R}^{N_{p}\times d_{t}}$ denote the learnable prompt from client $k$, where $N_{t}$ and $N_{p}$ are the lengths of the fixed and learnable prompts, $d_{t}$ is the dimension of hidden embedding, respectively. The full prompt can be formed as $P_{c}^{(k)}=[\delta_{k};\mathbf{p}_{c}]$. The text feature for class $c$ is computed by feeding the prompt into a text encoder $\mathcal{T}(\cdot)$: $\mathbf{f}_{t}^{c}=\mathcal{T}(P_{c}^{(k)})$. Given an input image $\mathbf{x}$, the image feature is extracted by the frozen image encoder $f$: $\mathbf{f}_{v}=f(\mathbf{x})$. The similarity score between the image and class-$c$ textual features is calculated as $\rho_{c}=\operatorname{Cosine}(\mathbf{f}_{t}^{c},\mathbf{f}_{v})$, and the prediction probability is computed via softmax over similarities: $p(\hat{y}=c\mid\mathbf{x})=\frac{\exp(\rho_{c}/\tau)}{\sum_{j=1}^{K}\exp(\rho_{j}/\tau)}$. The training objective is the cross-entropy loss between the similarity-based prediction and the ground-truth label: $\mathcal{L}_{\text{ce}}=\ell(\boldsymbol{\rho},\mathbf{e}_{y})$, where $\boldsymbol{\rho}=[\rho_{1},\dots,\rho_{K}]$, $\mathbf{e}_{y}$ is the one-hot label vector, and $\ell$ denotes the cross-entropy loss function. After local training, client $C_{k}$ uploads its learned prompt $\delta_{k}$ to the server. The server then aggregates the local prompts $\{\delta_{k}\}_{k=1}^{K}$ to construct a global prompt $\delta_{g}$, typically by weighted averaging: $\delta_{g}=\sum_{k=1}^{K}p_{k}\cdot\delta_{k}$.
4 The Proposed GPR-NIAM Method
In this section, we propose the non-interfering attention masking method for one-shot federated prompt learning, called GPR-NIAM. Its core idea is to preserve the transferable knowledge by reducing the influence of learnable prompt tokens on the input text tokens, as illustrated in Figure 2.
4.1 The Attention Isolation (AttnIso) Module
Prompt tuning typically focuses on fitting local training tasks. However, data heterogeneity often leads to divergent optimization across clients and weakens its cross-task generalization. To address this, the AttnIso module designs attention masks to selectively block the information flow from learnable prompt tokens to original text tokens (e.g. “a photo of a [CLASS]”). We denote the text attention matrix as $M\in\mathbb{R}^{1+N_{p}+N_{t}}$, where $N_{p}$ is the number of learnable prompts, $N_{t}$ is the length of input textual feature tokens. The additional 1 corresponds to the [EOS] token. In addition, we apply customized attention control for the [EOS] token to ensure proper aggregation of contextual information from both the prompt and the input tokens.
To achieve the above goal, the text attention matrix $M$ can be formulated as follows,
$M^{i,j}=\begin{cases}-\infty,&\begin{aligned} &\text{if }i\in[1,N_{p}]\land j\in[N_{p}+1,N_{p}+N_{t}]\quad&\\ &\text{or }i\in[N_{p}+1,N_{p}+N_{t}]\land j\in[1,N_{p}]\quad&\\ &\text{or }j=N_{p}+N_{t}+1\quad&\end{aligned}\quad\text{(Hard Masking)}\\ \lambda,&\text{if }i=N_{p}+N_{t}+1\land j\in[N_{p}+1,N_{p}+N_{t}]\quad\text{(Reweighting)}\\ 0,&\text{otherwise}\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{(Full Attention)}\end{cases}$ (1)
where $M^{i,j}$ denotes the $i$-th row and $j$-th column element of $M$, and $\land$ denotes logical AND. Moreover, to integrate the attention mask $M$ into the text encoder $\mathcal{T}$, we replace the original computation with the following:
$\begin{aligned} \mathbf{y}^{(l+1)}=\mathcal{T}_{l+1}(\mathbf{y}^{(l)},&M)=\text{Concat}(h_{1},h_{2},\dots,h_{H})W^{O}=[\tilde{p}_{l+1}^{1},\dots,\tilde{p}_{l+1}^{N_{p}},\tilde{t}_{l+1}^{1},\dots,\tilde{t}_{l+1}^{N_{t}},\tilde{s}_{l+1}^{\text{eos}}],\\ &\text{where }h_{i}=\operatorname{softmax}\left(\frac{Q_{i}K_{i}^{T}}{\sqrt{d_{k}}}+M\right)V_{i},\quad i=1,\dots,H\end{aligned}$
where $\mathbf{y}^{(l)}\in\mathbb{R}^{(N_{p}+N_{t}+1)\times d_{t}}$ and $\mathbf{y}^{(l+1)}\in\mathbb{R}^{(N_{p}+N_{t}+1)\times d_{t}}$ are the input and output of the $(l+1)$-th encoder layer. $\tilde{p}_{l+1}^{i}$, $\tilde{t}_{l+1}^{j}$, and $\tilde{s}_{l+1}^{\text{eos}}$ represent the output embeddings of the $i$-th prompt token, the $j$-th text token, and the [EOS] token, respectively. $Q_{i},K_{i},V_{i}\in\mathbb{R}^{(N_{p}+N_{t}+1)\times d_{k}}$ are the query, key, and value matrices for the $i$-th attention head, $d_{k}=d_{t}/H$ denotes the hidden dimension of each attention head. $\text{Concat}(h_{1},\dots,h_{H})\in\mathbb{R}^{(N_{p}+N_{t}+1)\times Hd_{k}}$ denotes the concatenated outputs of all heads, and $W^{O}\in\mathbb{R}^{Hd_{k}\times d_{t}}$ is the output projection matrix mapping the result back to $d_{t}$. Following the masked encoding process, we obtain the final output of the text encoder as: $\mathbf{y}^{(L)}=[\tilde{p}_{L}^{1},\dots,\tilde{p}_{L}^{N_{p}},\tilde{t}_{L}^{1},\dots,\tilde{t}_{L}^{N_{t}},\tilde{s}_{L}^{\text{eos}}]$. Afterwards, $\tilde{s}_{L}^{\text{eos}}$ is projected into a shared embedding space via a pre-trained head $H_{t}(\cdot)$. On the visual side, the input image $\mathbf{x}$ is encoded by a frozen vision encoder $f(\cdot)$ to obtain the visual representation. Overall, it focuses on minimizing the objective:
$\min_{\delta_{1},\dots,\delta_{K}}\sum_{k=1}^{K}\frac{|\mathcal{D}_{k}|}{|\mathcal{D}|}\cdot\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{k}}\left[-\log\left(\frac{\exp\left(\operatorname{sim}\left(f_{v}(\mathbf{x}),H_{t}\left(\tilde{s}^{(y)}_{L,k}\right)\right)\right)}{\sum_{c=1}^{C}\exp\left(\operatorname{sim}\left(f_{v}(\mathbf{x}),H_{t}\left(\tilde{s}^{(c)}_{L,k}\right)\right)\right)}\right)\right$],
where $\tilde{s}^{(c)}_{L,k}$ denotes client $k$’s [EOS] embedding for class $c$, $\operatorname{sim}(\cdot,\cdot)$ denotes cosine similarity function. The learnable prompts can be optimized by gradient descent in all clients:
$\delta_{k}\leftarrow\delta_{k}-\eta\cdot\frac{1}{|\mathcal{B}_{k}|}\sum_{(\mathbf{x},y)\in\mathcal{B}_{k}}\nabla_{\delta_{k}}\left[-\log\left(\frac{\exp\left(\operatorname{sim}\left(f_{v}(\mathbf{x}),H_{t}\left(\tilde{s}^{(y)}_{L,k}\right)\right)\right)}{\sum_{c=1}^{C}\exp\left(\operatorname{sim}\left(f_{v}(\mathbf{x}),H_{t}\left(\tilde{s}^{(c)}_{L,k}\right)\right)\right)}\right)\right],\quad\forall k\in[K],$
where $\eta$ denotes the learning rate, $\mathcal{B}_{k}$ is a batch data.
Figure 2: Illustration of the GPR-NIAM framework. The Attention Isolation module preserves the original text embeddings to reduce the forgetting of generalized knowledge. The global prompt is refined using visual prototypes from all clients to mitigate data heterogeneity.
4.2 Cross-Silo Collaborative Refinement (CSCR) Module
To further alleviate the inconsistency caused by data heterogeneity, the CSCR module calibrates the global prompt through prototype-guided refinement across clients. To support this, each client $C_{k}$ extracts class-wise visual prototypes from its local dataset $\mathcal{D}_{k}$. For class $c$, the prototype is computed as a randomly weighted average of the feature vectors from local samples in that class:
$f_{v,k}^{c}=\sum_{\mathbf{x}\in\mathcal{D}_{k}^{c}}\alpha_{\mathbf{x}}\cdot f_{v}(\mathbf{x}),\quad\text{where }\sum_{\mathbf{x}\in\mathcal{D}_{k}^{c}}\alpha_{\mathbf{x}}=1,\;\alpha_{\mathbf{x}}\sim\mathcal{U}(0,1).$ (5)
where $\alpha_{\mathbf{x}}$ denotes a random weight for each sample $\mathbf{x}$ in class $c$, drawn from a uniform distribution and normalized to ensure the weights sum to 1. Notably, we can randomly sample the weight vector $n$ times to generate multiple diverse prototypes per class $\mathbf{P}_{k}^{c}=\left[f_{v,k}^{c,1},f_{v,k}^{c,2},\dots,f_{v,k}^{c,n}\right]$, which enhances representational diversity, which differs from existing methods qi2023cross ; meng2024improving ; qi2024cross ; meng2025robust ; qi2025cross . $f_{v,k}^{c,i}$ denotes $i$-th visual prototypes of class $c$ in client $k$. Afterwards, each client $C_{k}$ sends $\mathbf{P}_{k}=\{\mathbf{P}_{k}^{1},\dots,\mathbf{P}_{k}^{N}\}$ to the server, which aggregates all received prototypes into a global pool $\mathcal{P}$:
$\mathcal{P}=\bigcup_{k=1}^{K}\mathbf{P}_{k}=\left\{\mathbf{P}_{k}^{c}\mid k\in[K],\;c\in[n]\right\},$ (6)
To calibrate the global prompt $\delta_{g}$, we leverage the global prototype pool $\mathcal{P}=\{\mathcal{P}^{1},\dots,\mathcal{P}^{C}\}$ and enforce the alignment between textual representations and visual prototypes via a class-level contrastive objective. Specifically, we design the following loss function:
$\min_{\delta_{g}}-\frac{1}{|\mathcal{P}|}\sum_{f\in\mathcal{P}}\log\sigma\left(\operatorname{sim}\left(f,\;H_{t}(\mathcal{T}(\delta_{g},\text{t}_{c}))\right)-\log\sum_{j=1}^{C}\exp\left(\operatorname{sim}\left(f,\;H_{t}(\mathcal{T}(\delta_{g},\text{t}_{j}))\right)\right)\right),$
where $\sigma(\cdot)$ is the sigmoid function, $\text{t}_{c}$ denotes ’a photo of c-th classname’. We then follow the gradient descent method to obtain the optimal parameters:
$\delta_{g}\leftarrow\delta_{g}-\eta\cdot\frac{1}{|\mathcal{P}|}\sum_{(f,c)\in\mathcal{P}}\nabla_{\delta_{g}}\log\sigma\left(\operatorname{sim}\left(f,\;H_{t}(\mathcal{T}(\delta_{g},\text{t}_{c}))\right)-\log\sum_{j=1}^{C}\exp\left(\operatorname{sim}\left(f,\;H_{t}(\mathcal{T}(\delta_{g},\text{t}_{j}))\right)\right)\right),$
where the global prompt $\delta_{g}$ is initialized as the weighted average of all client-specific prompts (i.e., $\delta_{g}\leftarrow\sum_{k=1}^{K}\frac{|\mathcal{D}_{k}|}{|\mathcal{D}|}\cdot\delta_{k}$), and $\eta$ is a learning rate.
4.3 Training Procedure
In this section, we provide the detailed training procedure of GPR-NIAM in Algorithm 1. As a two-stage optimization strategy, GPR-NIAM consists of a local prompt tuning phase and a global prompt refinement phase. Notably, it can be easily integrated into multiple prompt learning methods such as CoOp and TCP. In the first stage, each client initializes its local prompt from the global prompt and updates it on private data with masked attention learning. Simultaneously, class-wise visual prototypes are obtained by performing randomly weighted aggregation over local feature representations. Subsequently, the global prompt is optimized to align with all class prototypes in this pool via cross-silo collaborative retraining. With the two-stage design, GPR-NIAM alleviates the challenges posed by heterogeneous client data and promotes the retention of transferable knowledge for unseen tasks.
Moreover, the local prompt tuning stage is optional. Firstly, this stage provides a more meaningful initialization for the global prompt by adapting it to all clients data. Secondly, it generates class-wise visual prototypes to guide global refinement. While the framework can still perform global refinement without this phase, since it does not affect prototype modeling, omitting it might result in less effective initialization and slightly limit the overall refinement performance.
Algorithm 1 GPR-NIAM
1:Initialize the global prompt parameter $\delta^{0}$
2:// Local Prompt Tuning Stage
3:for each client $C_{k}\in\mathcal{C}$ in parallel do
4:  Initialize local prompt parameter $\delta_{k}=\delta^{0}$
5:  for $e=1,\dots,E$ do
6:   Sample a mini-batch $\zeta$ from local data $\mathcal{D}_{k}$
7:   Update local prompt $\delta_{k}$ using masked attention by Eq. (4)
8:  end for
9:  Compute visual prototype features for each class using Eq. (5)
10:end for
11:// Global Prompt Refinement Stage
12: Aggregate all visual prototypes to form a global prototype pool $\mathcal{P}$ by Eq. (6)
13:for $e^{\prime}=1,\dots,E^{{}^{\prime}}$ do
14:  Sample a mini-batch of prototypes $\zeta^{{}^{\prime}}$ from $\mathcal{P}$
15:  Update global prompt $\delta_{g}$ via prototype-guided alignment using Eq. (8)
16:end for
5 Experimental Evaluation
5.1 Experiment Settings
Datasets
Following existing works pan2024federated ; qiu2024federated , experiments are conducted on CIFAR10 krizhevsky2009learning , OxfordPets parkhi2012cats , Caltech101 fei2004learning , DTD cimpoi2014describing , FGVCAircraft maji2013fine , Flowers102 nilsback2008automated , StanfordCars krause20133d , and UCF101 soomro2012ucf101 to evaluate base-to-base/novel generalization, and on Office-Home torralba2011unbiased and DomainNet peng2019moment to assess leave-one-domain-out generalization. Their statistics are shown in Table 4.
Evaluation Metrics
Following studies qi2023cross ; qi2024cross ; liang2025tta , we report the Top-1 Accuracy (Acc) results. For base-to-base/novel generalization tasks, we also compute the harmonic mean ($\text{HM}=\frac{2\cdot\text{Acc}_{\text{base}}\cdot\text{Acc}_{\text{novel}}}{\text{Acc}_{\text{base}}+\text{Acc}_{\text{novel}}}$) of accuracy on base ($\text{Acc}_{\text{base}}$) and novel classes ($\text{Acc}_{\text{novel}}$) to evaluate overall performance.
Table 1: Performance comparison between GPR-NIAM and baselines on 8 datasets with $\beta=0.5$ in the base-to-base/novel generalization cases.
(a) Average over 8 datasets (b) CIFAR10 (c) OxfordPets
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 65.79 74.32 69.70 CLIP 95.30 95.50 95.39 CLIP 81.53 95.33 87.89
PromptFL 72.57 72.80 72.64 PromptFL 95.14 96.26 95.69 PromptFL 91.31 93.70 92.48
RPO 70.52 73.63 71.99 RPO 95.68 94.74 95.20 RPO 89.00 95.49 92.13
FedTPG 70.46 73.09 71.69 FedTPG 95.38 96.56 95.96 FedTPG 91.66 95.13 93.36
TCP 72.46 74.60 73.41 TCP 95.68 96.52 96.09 TCP 90.01 95.39 92.62
PromptFolio 73.25 73.35 73.15 PromptFolio 96.04 96.16 96.06 PromptFolio 91.13 96.10 93.54
DP-FPL 71.90 71.75 69.88 DP-FPL 95.42 95.18 95.29 DP-FPL 91.97 94.80 93.36
FedDISC 56.43 54.84 53.87 FedDISC 94.30 92.82 93.55 FedDISC 83.83 72.07 77.50
FedELMY 76.62 48.23 60.03 FedELMY 95.32 40.00 79.66 FedELMY 89.28 76.47 82.38
GPR-NIAMP 75.56 73.28 74.24 GPR-NIAMP 95.66 97.00 96.32 GPR-NIAMP 91.18 95.89 93.47
GPR-NIAMT 76.43 74.48 75.30 GPR-NIAMT 95.84 96.48 96.15 GPR-NIAMT 92.48 95.14 93.79
(d) Caltech101 (e) DTD (f) FGVCAircraft
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 82.27 94.05 87.77 CLIP 53.35 59.05 56.06 CLIP 26.51 31.21 28.67
PromptFL 88.27 93.88 90.99 PromptFL 60.06 53.86 56.79 PromptFL 31.19 30.91 31.05
RPO 89.90 93.80 91.81 RPO 54.62 59.17 56.81 RPO 29.27 30.43 29.84
FedTPG 89.12 92.61 90.83 FedTPG 56.71 55.43 56.06 FedTPG 29.57 30.25 29.90
TCP 88.20 95.33 91.62 TCP 61.68 55.48 58.41 TCP 30.77 32.71 31.71
PromptFolio 87.28 93.12 90.10 PromptFolio 61.11 54.58 57.66 PromptFolio 31.37 32.39 31.87
DP-FPL 87.92 93.46 90.60 DP-FPL 62.26 49.63 55.24 DP-FPL 28.85 28.15 28.49
FedDISC 44.75 85.22 58.68 FedDISC 71.52 38.28 49.87 FedDISC 20.87 13.20 16.17
FedELMY 87.71 89.21 88.45 FedELMY 70.13 16.18 26.29 FedELMY 33.77 19.48 24.70
GPR-NIAMP 91.59 94.22 92.89 GPR-NIAMP 68.40 53.50 60.04 GPR-NIAMP 30.95 29.41 30.16
GPR-NIAMT 92.65 94.73 93.68 GPR-NIAMT 69.90 55.91 62.13 GPR-NIAMT 32.15 31.12 31.62
(g) Flowers102 (h) StanfordCars (i) UCF101
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 62.39 78.10 69.36 CLIP 56.87 69.40 62.51 CLIP 68.12 71.92 69.97
PromptFL 79.88 75.15 77.44 PromptFL 62.96 67.30 65.02 PromptFL 71.82 71.55 71.68
RPO 74.92 76.21 75.56 RPO 58.64 68.73 63.29 RPO 72.19 70.49 71.33
FedTPG 72.30 77.26 74.70 FedTPG 58.67 69.78 63.74 FedTPG 70.34 67.74 69.01
TCP 77.25 78.73 77.98 TCP 60.19 71.18 65.22 TCP 75.93 71.52 73.65
PromptFolio 81.04 74.51 77.63 PromptFolio 61.22 69.91 65.27 PromptFolio 76.88 69.75 73.14
DP-FPL 74.63 77.26 75.92 DP-FPL 59.07 66.05 62.36 DP-FPL 75.14 69.49 72.20
FedDISC 24.95 32.21 28.11 FedDISC 42.77 49.92 46.07 FedDISC 68.49 55.03 61.03
FedELMY 89.79 55.15 68.33 FedELMY 67.09 47.97 55.94 FedELMY 79.94 41.41 54.56
GPR-NIAMP 84.54 78.52 81.42 GPR-NIAMP 63.59 67.48 65.48 GPR-NIAMP 78.57 70.28 74.19
GPR-NIAMT 86.88 77.05 81.67 GPR-NIAMT 63.54 69.18 66.24 GPR-NIAMT 78.04 76.27 77.14
Implementation Details
In all experiments, we set local training and global refinement epochs to 10, using the SGD optimizer with a learning rate of 0.001, a weight decay of 0.001, and a batch size of 32. The number of communication rounds is 1. We simulate 50 clients for CIFAR-10 and 10 clients for other datasets in base-to-base/novel generalization, while using 9 clients for Office-Home and 15 for DomainNet. To simulate non-IID data, we adopt a Dirichlet distribution with parameter $\beta=0.5$. Each client generates $n\in\{5,10,20\}$ visual prototypes per class, and the reweighting parameter $\lambda$ is selected from $\{0.2,0.5,0.7,1.0\}$. More details and results can be found in the Appendix. And each client has one NVIDIA RTX 3090 with 24 GB GPU for training.
5.2 Results and Discussion
This section compares GPR-NIAM-enhanced variants (GPR-NIAMP and GPR-NIAMT, derived by integrating GPR-NIAM into PromptFL and TCP, respectively) with nine baselines: Zero-Shot CLIP radford2021learning with hand-crafted text prompt template, e.g., “a photo of a [class]”, PromptFL guo2023PromptFL , RPO lee2023read , FedTPG qiu2024federated , TCP yao2024tcp , PromptFolio pan2024federated , DP-FPL tran2025privacy , FedDISC yang2024exploring , FedELMY wang2024one . For fair comparison, all methods use the same backbone: a frozen CLIP with ViT-B16 radford2021learning . Following prior work, the learnable prompt is a $10\times 512$ parameter matrix, with 10 prompt tokens and a 512-dimensional embedding.
Base-to-Base/Novel Generalization. We evaluate the generalization capability of GPR-NIAM by comparing its performance on base and novel classes, and the overall harmonic mean (HM) across all benchmarks. The results are presented in Table 1 and key observations are summarized as follows:
  • GPR-NIAM attains second-best performance on base classes among all methods, showing particular competitiveness against FedMLEY while maintaining more balanced capabilities across all classes.
  • GPR-NIAM yields competitive performance on novel classes. Notably, it ranks first on Flowers102 and UCF101, underscoring its strong generalization capability to unseen tasks. Moreover, FedDISC may fail on both base and novel classes due to inevitable gaps between synthetic and real samples.
  • Benefiting from the synergy of attention masking and cross-silo collaborative retraining, GPR-NIAM achieves the best HM scores across all benchmarks.
Table 2: Performance comparison on Office-Home and DomainNet under the leave-one-domain-out setting, with each domain’s data randomly split across 3 clients.
Method Office-Home DomainNet
Art Clipart Product RealWorld Average Clipart Infograph Painting Quickdraw Real Sketch Average
CLIP 74.57 62.33 80.17 79.38 74.08 92.87 77.02 85.35 43.06 98.61 94.22 81.85
PromptFL 74.24 66.20 82.47 79.77 75.17 93.87 80.82 87.76 49.47 98.72 94.82 84.24
RPO 74.86 64.33 84.31 80.74 76.06 93.53 78.32 86.15 49.35 98.83 94.87 83.50
FedTPG 74.04 64.55 83.32 80.01 75.48 95.14 80.93 88.25 52.89 98.97 95.40 85.26
TCP 75.31 66.20 83.17 80.97 76.41 94.13 79.80 88.05 45.85 98.80 95.34 83.66
PromptFolio 77.21 67.74 84.74 81.98 77.92 94.87 81.19 88.61 47.59 98.85 95.02 84.35
DP-FPL 75.03 65.91 83.71 79.82 76.12 95.03 78.41 86.74 54.03 98.83 94.85 84.64
FedDISC 72.27 65.40 81.77 79.11 74.63 93.82 80.42 86.17 43.74 98.66 94.05 82.81
FedELMY 75.11 66.34 84.12 79.69 76.31 94.21 77.65 86.81 48.78 98.83 94.88 83.53
GPR-NIAMP 78.32 71.40 87.61 84.96 80.57 94.72 80.13 88.70 56.98 99.18 94.88 85.76
GPR-NIAMT 79.39 72.06 88.03 85.49 81.24 95.03 80.82 87.92 56.32 99.08 95.60 85.79
Leave-One-Domain-Out Generalization. Table 2 presents the domain generalization results under a leave-one-domain-out setting. The main findings are summarized below:
  • GPR-NIAM consistently achieve the highest average accuracy across most domains. Notably, GPR-NIAMT ranks first in 9 out of 12 domains, highlighting its strength in cross-domain generalization, especially in challenging cases like Infograph and Quickdraw.
  • By combining token-level and client-level perspectives, GPR-NIAM outperforms single-view methods such as PromptFL and FedTPG, which benefits from attention masking for preserving original semantics and a unified global prompt for cross-domain knowledge integration.
5.3 Ablation Study
This section presents an ablation study to examine the effects of key components in GPR-NIAM, including the Attention Isolation (AttnIso) module and the Cross-Silo Collaborative Retraining (CSCR) module. The results are reported in Table 3.
[…]
  • The CSCR module plays a vital role in optimizing base-class performance by learning a unified prompt that mitigates inter-client discrepancies, which can alleviate the negative impact of data heterogeneity. However, its contribution to novel-class generalization remains limited due to insufficient adaptability to unseen categories.
  • By integrating both modules, GPR-NIAM achieves an optimal balance between acquiring new knowledge and preserving generalization. Importantly, this also suggests a promising direction for future work by combining complementary strategies rather than relying on a single one.
Table 3: Ablation results averaged over 8 datasets (base-to-base/novel generalization) and across all domains (leave-one-domain-out generalization).
Method Base to Base/Novel Generalization Leave-One-Domain-Out Generalization
Base Novel HM Office-Home DomainNet
PromptFL 64.06 71.48 66.98 75.17 84.24
[…]
+ CSCR 73.34 64.89 67.24 78.59 85.11
[…]
Figure 3: The impact of hyperparameters on performance is evaluated by varying $\lambda$ from $\{0,0.2,0.5,0.7,1.0\}$, the number of prototypes per class $n$ from $\{3,5,10\}$ on three datasets.
5.4 Sensitivity Analysis
This section evaluates the robustness of GPR-NIAM across varying hyperparameter settings on the DTD, Flowers102 and Office-Home datasets. Specifically, we investigate the effects of hyperparameters $n$ (the number of prototypes generated per class on the client side) and $\lambda$ (the weighting factor in the AttnIso module), evaluated over $\{3,5,10\}$ and $\{0,0.2,0.5,0.7,1.0\}$, respectively. As depicted in Figure 3, GPR-NIAM consistently achieves higher overall HM scores than its baselines in all cases of the base-to-base/novel generalization task, showcasing its superior balance between retaining transferable knowledge and learning task-specific knowledge. On the one hand, GPR-NIAM demonstrates consistent improvements over baselines on base classes, which benefits from the CSCR module’s capacity to alleviate data heterogeneity. On the other hand GPR-NIAM exhibits competitive performance on novel-class generalization, frequently outperforming other methods, which is attributed to the attention masking mechanism of the AttnIso module. Furthermore, in the Leave-One-Domain-Out Generalization task, similar conclusions are drawn, as GPR-NIAM outperforms the baseline across multiple domains.
5.5 Effectiveness of Global Prompt Refinement
Figure 4: Visualization of similarity heatmaps between text features and image patches before and after prompt refinement on Caltech101. In the similarity maps, red and yellow indicate higher similarity between image patches and the text features, while blue represents lower similarity.
This section aims to evaluate the effectiveness of the global prompt refinement method. Specifically, we visualize the similarity maps between text features and image patches before and after refinement, as shown in Figure 4. Intuitively, the similarity responses between visual targets and textual semantics are often scattered and misaligned before refinement. For example, in the “Dollar Bill” and “Bonsai” samples, the high similarity responses in the second row primarily concentrate on background regions near the image boundaries. In contrast, the responses become significantly more concentrated on semantically relevant regions, such as the face on the bill or the foliage of the bonsai after refinement. Notably, this advantage is also evident in novel classes. For instance, the refined prompts enable the model to better localize foreground objects while suppressing irrelevant background activations in “Kangaroo” and “Nautilus” samples. Moreover, it becomes tightly aligned with the spiral structure of the shell in the “Nautilus”, indicating enhanced semantic grounding.
Figure 5: Comparison between Random and Local Prompt-Tuning for Initialization.
Figure 6: Failure case analysis. Most GPR-NIAM errors overlap with zero-shot CLIP.
5.6 Evaluation of Local Prompt Tuning as Initialization
This section evaluates the performance of random initialization and local prompt tuning (LPT) on Flowers102, UCF101 and DTD datasets. As shown in Figure 6, LPT yields superior performance compared to random initialization on all cases. This can be attributed to the local training that can be more aligned with the target data, allowing the model to better capture domain-specific features. In particular, LPT outperforms the random initialization by 6% on the DTD dataset for base classes, which highlights the method’s ability to adapt more efficiently to the nuances of the texture.
5.7 Failure Case Analysis
This section reports the percentage of cases where the zero-shot CLIP model also failed in instances where the GPR-NIAM model made incorrect predictions (Figure 6). Clearly, there is a high overlap in the failure cases of both models. This reflects that the imprecision of zero-shot generalized knowledge hinders the model’s generalization, which also provides insights for future work to address the challenges posed by the lack of domain-specific adaptation in generalization models.
6 Conclusions and Future Work
In this paper, we balance the model capability to learn task-specific knowledge for downstream tasks with its capacity to retain generalized knowledge. Specifically, we propose GPR-NIAM for one-shot FPL, which introduces an attention masking mechanism to limit the interaction between the learnable prompt tokens and the original text tokens. Furthermore, it learns a unified prompt through global prompt refinement, thereby mitigating the data heterogeneity issue. Experimental results demonstrate that GPR-NIAM significantly improves the performance of global prompts across diverse visual tasks.
Despite the promising results of GPR-NIAM, there remains room for improvement. The communication overhead from prototype transmission may become a bottleneck in large-scale federated systems with limited bandwidth, which we plan to address in future work. And both the multi-label problem kou2025labeldistributionlearningbiased ; ijcai2024p478 and real-world medical imaging tasks feng2023unsupervised ; feng2025neighbor become considerably more challenging in federated settings. In addition, several issues discussed in Section 5.7 remain open and deserve further investigation.
Appendix A Appendix
A.1 Datasets
To conduct our evaluation, we selected 10 diverse visual classification datasets as benchmarks. Table 4 provides a detailed overview, including the number of classes, the size of training and testing sets, the number of domains, and the corresponding federated settings.
For datasets with multiple domains, we adopt the commonly used Office-Home benchmark, which includes four domains: Art, Clipart, Product, and RealWorld. In addition, we utilize DomainNet, a large-scale dataset spanning six domains: Clipart, Infograph, Painting, Quickdraw, Real, and Sketch. We focus on training with a selected subset of 20 classes from each dataset. Figure 7 illustrates representative raw instances from the two multi-domain datasets.
Table 4: Statistics and federated settings of the datasets used in our experiments.
Dataset Classes Train Test Domains Federated Settings
Base Novel Clients Heterogeneity
CIFAR10 10 25,000 5,000 5,000 1 5/10/20 0.1/0.5
OxfordPets 37 1,785 1,782 1,887 1 5/10/20 0.1/0.5
Caltech101 100 3,310 1,416 1,178 1 5/10/20 0.1/0.5
DTD 47 1,440 864 828 1 5/10/20 0.1/0.5
FGVCAircraft 100 3,333 1,667 1,666 1 5/10/20 0.1/0.5
Flowers102 102 2,807 343 475 1 5/10/20 0.1/0.5
StanfordCars 196 4,052 4,002 3,998 1 5/10/20 0.1/0.5
UCF101 101 3,926 1,895 1,888 1 5/10/20 0.1/0.5
DomainNet 20 30,777 6 9/15
Office-Home 65 15,588 4 15/25
Figure 7: Illustrative examples of class instances across domains: “Ambulance” from DomainNet and “Bed” from Office-Home.
A.2 Experimental Results
A.2.1 Base-to-Base/Novel Generalization under High-Level Data Heterogeneity
Table 5: Performance comparison between GPR-NIAM and baselines on 8 datasets with $\beta=0.1$ in the base-to-base/novel generalization setting. The best performance is indicated in bold.
(a) Average over 8 datasets (b) CIFAR10 (c) OxfordPets
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 65.79 74.32 69.70 CLIP 95.30 95.50 95.39 CLIP 81.53 95.33 87.89
PromptFL 69.15 72.50 70.71 PromptFL 94.72 95.68 95.19 PromptFL 89.33 95.91 92.51
RPO 67.81 73.66 70.57 RPO 95.10 94.76 94.92 RPO 86.19 95.54 90.63
FedTPG 67.06 71.71 69.20 FedTPG 94.30 96.64 95.45 FedTPG 86.58 94.70 90.46
TCP 70.30 74.19 72.13 TCP 95.30 95.66 95.47 TCP 90.01 96.23 93.02
PromptFolio 69.49 73.44 71.33 PromptFolio 95.20 96.74 95.96 PromptFolio 90.90 95.17 92.99
DP-FPL 71.90 71.75 69.88 DP-FPL 94.24 95.42 94.82 DP-FPL 91.30 95.12 93.17
FedDISC 52.27 54.40 50.47 FedDISC 92.47 91.37 91.91 FedDISC 81.48 66.74 73.38
FedELMY 68.05 47.29 53.41 FedELMY 89.84 28.90 43.73 FedELMY 86.70 82.77 84.69
GPR-NIAMP 73.87 72.84 73.18 GPR-NIAMP 95.88 96.84 96.35 GPR-NIAMP 90.57 95.23 92.84
GPR-NIAMT 73.62 74.48 73.97 GPR-NIAMT 96.03 96.24 96.13 GPR-NIAMT 91.86 94.38 93.10
(d) Caltech101 (e) DTD (f) FGVCAircraft
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 82.27 94.05 87.77 CLIP 53.35 59.05 56.06 CLIP 26.51 31.21 28.67
PromptFL 87.07 93.03 89.95 PromptFL 56.36 52.53 54.38 PromptFL 26.03 30.61 28.13
RPO 84.95 93.80 89.16 RPO 53.12 60.02 56.36 RPO 26.51 30.37 28.31
FedTPG 85.38 93.37 89.20 FedTPG 53.35 50.00 51.62 FedTPG 21.17 26.35 23.48
TCP 88.41 94.16 91.19 TCP 58.10 57.60 57.85 TCP 28.37 32.23 30.18
PromptFolio 82.41 92.61 87.21 PromptFolio 53.58 55.55 54.55 PromptFolio 28.13 34.15 30.85
DP-FPL 87.57 93.54 90.46 DP-FPL 51.85 45.65 48.55 DP-FPL 26.63 29.35 27.92
FedDISC 22.37 81.39 35.09 FedDISC 65.23 41.23 49.87 FedDISC 18.49 17.61 18.03
FedELMY 85.59 90.40 87.93 FedELMY 40.27 11.95 18.43 FedELMY 23.45 5.46 8.86
GPR-NIAMP 91.24 93.88 92.54 GPR-NIAMP 67.59 52.89 59.34 GPR-NIAMP 29.69 30.81 30.23
GPR-NIAMT 90.60 94.05 92.30 GPR-NIAMT 65.62 57.79 61.45 GPR-NIAMT 29.39 29.96 29.67
(g) Flowers102 (h) StanfordCars (i) UCF101
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 62.39 78.10 69.36 CLIP 56.87 69.40 62.51 CLIP 68.12 71.92 69.97
PromptFL 72.88 75.78 74.30 PromptFL 57.31 67.58 62.02 PromptFL 69.55 68.90 69.22
RPO 72.01 76.63 74.24 RPO 57.64 66.73 61.85 RPO 67.01 71.45 69.16
FedTPG 70.84 77.26 73.91 FedTPG 56.89 69.55 62.59 FedTPG 68.02 65.83 66.91
TCP 73.61 77.94 75.71 TCP 57.79 69.28 63.01 TCP 70.88 70.43 70.65
PromptFolio 75.51 76.42 75.96 PromptFolio 58.97 69.05 63.61 PromptFolio 71.29 67.84 69.52
DP-FPL 68.97 71.43 70.08 DP-FPL 53.99 63.65 58.43 DP-FPL 73.61 69.80 71.66
FedDISC 25.44 32.47 28.52 FedDISC 42.59 51.23 46.51 FedDISC 70.12 53.23 60.51
FedELMY 81.63 57.05 67.16 FedELMY 60.26 49.94 54.62 FedELMY 76.72 51.90 61.92
GPR-NIAMP 79.30 77.26 78.26 GPR-NIAMP 60.81 66.98 63.75 GPR-NIAMP 75.88 68.85 72.19
GPR-NIAMT 80.17 79.15 79.65 GPR-NIAMT 60.14 68.65 64.12 GPR-NIAMT 75.19 75.63 75.41
To more comprehensively evaluate the effectiveness of the methods, we test all approaches on eight datasets under a higher level of heterogeneity ($\beta=0.1$), focusing on their performance in the base-to-base/novel generalization setting. As presented in Table 5, our proposed method GPR-NIAM obtains the highest average performance on the eight datasets. Specifically, GPR-NIAMP and GPR-NIAMT achieve the top two performances on the base classes, ranking first and second, respectively. In terms of novel class generalization, GPR-NIAMT outperforms all baselines, demonstrating its superior capability to adapt to unseen categories. Furthermore, GPR-NIAMT also achieves the highest overall performance in terms of harmonic mean (HM), indicating a strong balance between knowledge retention and generalization. These results highlight the robustness and adaptability of the proposed method under highly heterogeneous federated settings.
A.2.2 Leave-One-Domain-Out Generalization with Increased Client Diversity
Table 6: Performance comparison between GPR-NIAM and baselines on the Office-Home and DomainNet datasets under the leave-one-domain-out generalization setting, where the data from each domain is randomly partitioned among 5 clients.
Method Office-Home DomainNet
Art Clipart Product RealWorld Average Clipart Infograph Painting Quickdraw Real Sketch Average
CLIP 74.57 62.33 80.17 79.38 74.08 92.87 77.02 85.35 43.06 98.61 94.22 81.85
PromptFL 73.71 64.49 82.81 80.39 75.35 94.93 80.09 87.94 52.68 99.00 95.63 85.04
RPO 73.83 62.13 79.95 79.36 73.81 93.66 77.31 84.65 44.46 98.83 93.87 82.13
FedTPG 74.90 65.54 82.85 80.37 75.91 94.35 79.76 87.40 48.05 98.92 95.89 84.06
TCP 75.77 65.10 83.10 80.99 76.24 93.92 79.72 87.67 46.82 98.96 95.00 83.68
PromptFolio 75.81 65.36 83.53 80.39 76.27 95.51 81.00 89.33 48.82 98.99 95.20 84.80
DP-FPL 74.82 63.55 82.85 78.79 75.00 94.77 80.20 89.00 48.35 98.78 94.82 84.32
FedDISC 71.19 62.48 81.31 78.35 73.48 94.23 80.25 83.85 44.21 98.39 92.49 82.23
FedELMY 75.45 64.29 82.36 77.58 74.92 94.15 75.48 86.89 47.39 98.72 95.13 82.96
GPR-NIAMP 78.16 72.76 89.09 83.68 80.92 94.45 80.60 87.50 58.68 99.10 94.91 85.87
GPR-NIAMT 77.66 71.95 89.47 85.93 81.25 95.19 80.49 87.23 57.96 99.07 95.37 85.88
To validate the effectiveness of the proposed GPR-NIAM under more challenging scenarios, we conduct experiments in a leave-one-domain-out generalization setting [] using the Office-Home and DomainNet datasets. In this setting, the data from each domain is randomly partitioned among 5 clients, and for each evaluation, one domain is held out entirely for testing while the remaining domains are used for federated training. As shown in Table 6, GPR-NIAMP and GPR-NIAMT consistently outperform existing baselines across both datasets. For the Office-Home, GPR-NIAMT achieves the highest average accuracy, with GPR-NIAMP closely following. Notably, GPR-NIAMT yields the best performance on the “RealWorld" and “Product" domains, while GPR-NIAMP leads on “Art" and “Clipart". For the DomainNet, both variants of GPR-NIAM also demonstrate strong cross-domain generalization, achieving top results in multiple domains. These results highlight the robustness of our method in handling both domain shift and non-IID data across clients. In summary, the proposed GPR-NIAM exhibits superior generalization performance under leave-one-domain-out conditions, demonstrating its effectiveness in federated scenarios with high domain and client-level variability.
A.2.3 Performance Comparison across Different Local Training Epoch Settings
Table 7: Performance comparison between GPR-NIAM and baselines on DTD and Flowers102 with local training epochs $\in\{1,5,20\}$ in the base-to-base/novel generalization setting.
(a) DTD (Epochs=1) (b) DTD (Epochs=5) (c) DTD (Epochs=20)
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 53.35 59.05 56.06 CLIP 53.35 59.05 56.06 CLIP 53.35 59.05 56.06
PromptFL 50.34 51.69 51.01 PromptFL 55.55 56.76 56.15 PromptFL 63.31 51.08 56.54
RPO 52.77 59.42 55.90 RPO 53.24 59.78 56.32 RPO 53.93 60.26 56.92
FedTPG 40.27 36.47 38.28 FedTPG 49.88 47.22 48.51 FedTPG 56.25 56.28 56.26
TCP 53.24 53.38 53.31 TCP 57.63 56.40 57.01 TCP 63.07 59.54 61.25
PromptFolio 52.54 53.50 53.02 PromptFolio 58.68 56.28 57.45 PromptFolio 58.68 57.60 58.13
DP-FPL 43.75 41.54 42.61 DP-FPL 59.25 50.12 54.30 DP-FPL 64.12 47.82 54.78
GPR-NIAMP 65.97 51.20 57.65 GPR-NIAMP 68.51 52.65 59.54 GPR-NIAMP 70.13 53.26 60.54
GPR-NIAMT 63.07 51.81 56.89 GPR-NIAMT 65.85 55.07 59.98 GPR-NIAMT 69.67 54.83 61.36
(d) Flowers102 (Epochs=1) (e) Flowers102 (Epochs=5) (f) Flowers102 (Epochs=20)
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 62.39 78.10 69.36 CLIP 62.39 78.10 69.36 CLIP 62.39 78.10 69.36
PromptFL 70.84 74.10 72.43 PromptFL 78.42 76.00 77.19 PromptFL 80.46 73.68 76.92
RPO 69.97 76.21 72.95 RPO 71.72 76.84 74.19 RPO 74.05 77.05 75.52
FedTPG 68.51 77.05 72.53 FedTPG 73.76 77.05 75.37 FedTPG 74.34 77.05 75.67
TCP 75.51 78.10 76.78 TCP 76.09 79.15 77.59 TCP 79.88 78.31 79.09
PromptFolio 73.76 76.42 75.06 PromptFolio 79.59 75.57 77.53 PromptFolio 83.09 74.52 78.57
DP-FPL 69.38 78.31 73.58 DP-FPL 73.76 79.57 76.55 DP-FPL 75.80 75.78 75.79
GPR-NIAMP 82.50 75.57 78.89 GPR-NIAMP 84.83 78.31 81.44 GPR-NIAMP 85.13 77.47 81.12
GPR-NIAMT 87.11 77.83 82.20 GPR-NIAMT 86.58 78.94 82.59 GPR-NIAMT 87.75 78.73 83.00
Table 8: Performance comparison between GPR-NIAM and baselines on StanfordCars and UCF101 with local training epochs $\in\{1,5,20\}$ in the base-to-base/novel generalization setting.
(g) StanfordCars (Epochs=1) (h) StanfordCars (Epochs=5) (i) StanfordCars (Epochs=20)
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 56.87 69.40 62.51 CLIP 56.87 69.40 62.51 CLIP 56.87 69.40 62.51
PromptFL 58.87 68.40 63.27 PromptFL 60.39 67.35 63.68 PromptFL 59.74 67.83 63.52
RPO 56.39 69.03 62.07 RPO 57.12 69.70 62.79 RPO 57.42 70.06 63.11
FedTPG 57.09 69.73 62.78 FedTPG 58.14 69.65 63.38 FedTPG 58.69 69.25 63.54
TCP 58.04 69.40 63.22 TCP 58.72 69.91 63.82 TCP 59.29 68.68 63.64
PromptFolio 58.69 69.83 63.78 PromptFolio 60.04 69.78 64.54 PromptFolio 61.46 69.10 65.06
DP-FPL 49.30 62.85 55.25 DP-FPL 59.02 68.33 63.33 DP-FPL 59.24 64.35 61.69
GPR-NIAMP 60.61 67.58 63.91 GPR-NIAMP 62.09 67.05 64.48 GPR-NIAMP 64.46 66.38 65.41
GPR-NIAMT 61.54 67.40 64.33 GPR-NIAMT 60.99 69.45 64.94 GPR-NIAMT 63.01 69.00 65.87
(j) UCF101 (Epochs=1) (k) UCF101 (Epochs=5) (l) UCF101 (Epochs=20)
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 68.12 71.92 69.97 CLIP 68.12 71.92 69.97 CLIP 68.12 71.92 69.97
PromptFL 69.55 69.01 69.28 PromptFL 73.03 72.08 72.55 PromptFL 68.07 64.24 66.10
RPO 65.96 70.55 68.17 RPO 66.80 71.50 69.07 RPO 68.75 71.13 69.92
FedTPG 64.69 64.08 64.39 FedTPG 68.86 67.47 68.16 FedTPG 68.86 66.73 67.77
TCP 73.82 72.47 73.13 TCP 75.83 72.27 74.00 TCP 76.09 73.00 74.51
PromptFolio 73.72 69.80 71.71 PromptFolio 76.56 70.12 73.20 PromptFolio 78.15 68.37 72.94
DP-FPL 72.13 70.65 71.38 DP-FPL 74.08 71.76 72.91 DP-FPL 73.24 68.16 70.61
GPR-NIAMP 76.88 68.53 72.47 GPR-NIAMP 79.10 68.75 73.56 GPR-NIAMP 78.89 69.38 73.83
GPR-NIAMT 77.78 74.62 76.17 GPR-NIAMT 78.15 75.79 76.95 GPR-NIAMT 78.04 76.58 77.31
To evaluate the effectiveness of the proposed GPR-NIAM under different levels of local training effort, this section investigates its performance across varying local training epochs. Specifically, we consider three settings with local training epochs $\in\{1,5,20\}$, under a data heterogeneity level ($\beta=0.5$) and with 10 clients participating in training. As shown in Table 7 and Table 8, both GPR-NIAMP{P} and GPR-NIAMT{T} consistently outperform the baselines across all datasets and training epochs. Notably, GPR-NIAMT{T} delivers the best overall performance in terms of harmonic mean (HM) in most settings, which demonstrates superior generalization across both base and novel classes. On Flowers102 and UCF101, GPR-NIAMT{T} shows strong robustness even when the number of local epochs increases. This highlights its ability to prevent overfitting to local distributions. Across datasets, both variants of GPR-NIAM remain stable and effective even as the local training dynamics change, outperforming methods that are more sensitive to overfitting or class imbalance under prolonged local updates.
A.2.4 Evaluating Method Scalability with Varying Client Numbers
Table 9: Performance comparison between GPR-NIAM and baselines on DTD and Flowers102 datasets under varying number of clients $K\in\{20,30\}$ in the base-to-base/novel generalization setting.
(a) DTD (K=20) (b) DTD (K=30)
Method Base Novel HM Method Base Novel HM
CLIP 53.35 59.05 56.06 CLIP 53.35 59.05 56.06
PromptFL 54.16 54.22 54.19 PromptFL 52.89 56.52 54.64
RPO 53.24 60.02 56.42 RPO 52.77 59.66 56.00
FedTPG 57.83 54.93 56.34 FedTPG 52.19 58.09 54.98
TCP 58.21 55.93 57.04 TCP 56.82 58.42 57.60
PromptFolio 58.18 52.91 55.41 PromptFolio 55.43 52.41 53.88
DP-FPL 56.13 50.72 53.29 DP-FPL 57.06 51.20 53.97
GPR-NIAMP 63.31 52.29 57.27 GPR-NIAMP 60.64 51.93 55.95
GPR-NIAMT 62.38 53.86 57.81 GPR-NIAMT 60.41 56.52 58.40
(d) Flowers102 (K=20) (e) Flowers102 (K=30)
Method Base Novel HM Method Base Novel HM
CLIP 62.39 78.10 69.36 CLIP 62.39 78.10 69.36
PromptFL 74.34 74.10 74.22 PromptFL 72.01 74.73 73.34
RPO 71.72 76.84 74.19 RPO 71.13 76.42 73.68
FedTPG 74.63 78.31 76.43 FedTPG 74.92 78.31 76.58
TCP 75.21 77.68 76.43 TCP 73.46 78.73 76.01
PromptFolio 75.51 78.31 76.88 PromptFolio 73.76 78.31 75.97
DP-FPL 74.34 74.10 74.22 DP-FPL 69.09 75.57 72.19
GPR-NIAMP 82.79 77.05 79.82 GPR-NIAMP 81.63 78.31 79.93
GPR-NIAMT 83.96 77.05 80.36 GPR-NIAMT 82.50 78.94 80.68
Table 10: Performance comparison between GPR-NIAM and baselines on StanfordCars and UCF101 datasets under varying number of clients $K\in\{20,30\}$ in the base-to-base/novel generalization setting.
(g) StanfordCars (K=20) (h) StanfordCars (K=30)
Method Base Novel HM Method Base Novel HM
CLIP 56.87 69.40 62.51 CLIP 56.87 69.40 62.51
PromptFL 59.82 67.03 63.22 PromptFL 54.42 66.33 59.79
RPO 57.22 69.65 62.83 RPO 57.24 69.60 62.82
FedTPG 59.24 66.73 62.76 FedTPG 59.42 69.08 63.88
TCP 58.77 67.78 62.95 TCP 59.32 68.21 63.45
PromptFolio 60.91 65.78 63.25 PromptFolio 59.59 65.25 62.29
DP-FPL 57.74 65.65 61.44 DP-FPL 58.37 67.75 62.71
GPR-NIAMP 60.04 68.23 63.87 GPR-NIAMP 60.09 67.40 63.54
GPR-NIAMT 61.02 69.40 64.94 GPR-NIAMT 60.14 69.38 64.43
(j) UCF101 (K=20) (k) UCF101 (K=30)
Method Base Novel HM Method Base Novel HM
CLIP 68.12 71.92 69.97 CLIP 68.12 71.92 69.97
PromptFL 75.67 68.75 72.04 PromptFL 71.34 66.04 68.59
RPO 66.80 71.34 69.00 RPO 66.75 71.39 68.99
FedTPG 74.88 74.78 74.83 FedTPG 76.09 74.62 75.35
TCP 74.08 73.83 73.96 TCP 73.50 75.58 74.53
PromptFolio 74.45 71.76 73.08 PromptFolio 77.09 74.04 75.54
DP-FPL 75.51 69.75 72.52 DP-FPL 72.71 68.96 70.79
GPR-NIAMP 76.41 69.01 72.52 GPR-NIAMP 76.62 72.50 74.50
GPR-NIAMT 77.04 76.21 76.62 GPR-NIAMT 77.62 76.05 76.83
This section aims to evaluate the scalability of the proposed GPR-NIAM with respect to the number of participating clients. We consider two settings with $K\in\{20,30\}$ under a heterogeneity level $\beta=0.5$ and fixed local training epochs of 5. As shown in Table 9 and Table 10, both GPR-NIAMP and GPR-NIAMT consistently outperform all baselines across datasets and client configurations. GPR-NIAMT achieves the best overall performance on most datasets and demonstrates strong generalization to novel classes. For example, on Flowers102 and UCF101, it achieves the highest harmonic mean (HM) under both client settings. Meanwhile, GPR-NIAMP maintains strong performance on base classes. These results show that GPR-NIAM maintains high performance and stability even as the number of clients increases, demonstrating its scalability and effectiveness in more challenging cases.
A.2.5 Evaluation of Module Contributions
Table 11: Ablation study on the effectiveness of NIAM and CSCR modules under the base-to-base/novel generalization. Results are reported on four datasets with PromptFL as the baseline.
(a) DTD (b) Flowers102
Method Base Novel HM Method Base Novel HM
PromptFL 55.55 56.76 56.15 PromptFL 78.42 76.00 77.19
+NIAM 55.78 54.73 55.25 +NIAM 76.13 77.78 76.94
+NIAM+CSCR 68.51 52.65 59.54 +NIAM+CSCR 84.83 78.31 81.44
(c) StanfordCars (d) UCF101
Method Base Novel HM Method Base Novel HM
PromptFL 60.39 67.35 63.68 PromptFL 73.03 72.08 72.55
+NIAM 55.17 68.13 60.96 +NIAM 67.96 73.02 70.39
+NIAM+CSCR 62.09 67.05 64.48 +NIAM+CSCR 79.10 68.75 73.56
We evaluate the contributions of the NIAM and CSCR modules by gradually adding them to the PromptFL baseline. Experiments are conducted on four datasets using 10 clients, with heterogeneity level $\beta=0.5$ and 5 local training epochs. As shown in Table 11, NIAM improves performance on novel classes by preserving generalized knowledge. CSCR enhances base class performance through visual supervision. Together, they deliver the best overall performance in terms of harmonic mean (HM), highlighting their complementary effects.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: This study outlines the research problem and its contributions in both the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations are discussed in the conclusion and future work sections.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not involve theoretical assumptions.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: This paper provides a detailed description of the experimental setup, including the range of parameter tuning, and the code will be uploaded as supplementary material.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code will be uploaded as supplementary material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental setup is clearly described in both the main experiment section and the appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: This paper adopts evaluation metrics and assessment methods consistent with existing studies.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The computational resources used are described in the experimental implementation details section.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in this paper fully conforms to the NeurIPS Code of Ethics in every respect.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: As discussed in the introduction, the study substantially reduces communication costs by avoiding multiple rounds of exchange, which is advantageous for deployment under stringent latency or bandwidth limitations.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper utilizes the fully open-source CLIP model and publicly available benchmark datasets that are widely used in existing literature. Therefore, no additional safeguards are necessary for responsible release.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: the paper cites all relevant works, and all utilized assets are open-source and used in accordance with their licenses.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The code developed for this work is included as anonymized supplementary material
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: This paper only uses LLMs for grammar checking and editing.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 1 Introduction
-Federated learning (FL) enables collaborative modeling across data sources in a privacy-preserving manner.
-It learns global knowledge by aggregating model weights or gradients instead of moving data to a central entity mcmahan2017communication ; cai2024lgfgad ; 9820618 ; yi2024federated ; yi2023pfedes ; 11091509 .
+Federated learning (FL) enables collaborative modeling across data sources in a privacy-preserving manner, which has gained traction across multiple domains qi2024attentive ; cai2024fgad ; guo2024sample ; li2020federated ; fu2025less .
+It learns global knowledge by aggregating model weights or gradients instead of moving data to a central entity cai2024lgfgad ; 9820618 ; yi2024federated ; yi2023pfedes .
 Recently, Federated Prompt Learning (FPL) has emerged to support parameter-efficient fine-tuning in this setting.
 It not only leverages pre-trained models (e.g., CLIP radford2021learning ) more effectively, but also reduces communication overhead between FL clients and the FL server by transmitting only learnable prompts qi2025federated ; zhao2023fedprompt ; ren2025advances ; fan2025ten .
-However, heterogeneous data distributions across clients often degrade model aggregation performance and slow down convergence yang2023efficient ; bai2024diprompt ; li2020federated ; guo2025federated ; guo2024sample .
-This is because significant differences in the update directions across local models can cause conflicts during aggregation pan2024federated ; cui2024harmonizing ; huang2024federated ; dong2022federated_FCIL ; fu2025federated .
+However, heterogeneous data distributions across clients often degrade model aggregation performance and slow down convergence yang2023efficient ; guo2025federated .
+This is because significant differences in the update directions across local models can cause conflicts during aggregation pan2024federated ; huang2024federated ; dong2022federated_FCIL ; fu2025federated ; zhang2024enabling .
 As a result, existing FPL methods still require multi-round communications between clients and the server to reconcile inconsistencies arising from client-specific training.
@@
 This poses a challenge in real-world scenarios with limited communication resources.
-To this end, one-shot federated learning (OSFL), which requires only a single round of client-server interaction, offers a promising direction for further exploration zhang2022dense ; zeng2024one ; diao2023towards ; yang2024exploring ; cai2024fgad .
+To this end, one-shot federated learning (OSFL), which requires only a single round of client-server interaction, offers a promising direction for further exploration zhang2022dense ; zeng2024one ; diao2023towards ; yang2024exploring ; 11091509 .
 OSFL significantly reduces the communication overhead by eliminating iterative exchanges, making it attractive for real-world deployments with strict latency or bandwidth constraints yao2025fedmho ; ma2025geometric ; hu2024fedmut .
@@
 Suppose there are $K$ clients denoted by $\mathcal{C}=\{C_{1},C_{2},\dots,C_{K}\}$, and each client $C_{k}\in\mathcal{C}$ holds a private dataset $\mathcal{D}_{k}$.
-The goal of traditional FL is to collaboratively train a global model $f$ by minimizing the objective: $\min_{f}\sum_{k=1}^{K}p_{k}\cdot\mathcal{L}(f,\mathcal{D}_{k})$, where $p_{k}=\frac{|\mathcal{D}_{k}|}{\sum_{j=1}^{K}|\mathcal{D}_{j}|}$ is the weight of client $k$, and $\mathcal{L}(\cdot)$ is a task-specific loss function.
+The goal of traditional FL is to collaboratively train a global model $f$ by minimizing the objective: $\min_{f}\sum_{k=1}^{K}p_{k}\cdot\mathcal{L}(f,\mathcal{D}_{k})$, where $p_{k}=\frac{|\mathcal{D}_{k}|}{\sum_{j=1}^{K}|\mathcal{D}_{j}|}$ is the weight of client $k$, and $\mathcal{L}(\cdot)$ is a task-specific loss function hu2024fedcross ; qi2025global ; li2025re ; zhangcausality .
 In FPL, each client $k$ optimizes a set of learnable prompts $\delta_{k}$ using its private dataset $\mathcal{D}_{k}$, while keeping the backbone model $f$ frozen: $\min_{\delta_{k}}\mathcal{L}(f;\delta_{k},\mathcal{D}_{k})$.
@@
 “a photo of a [CLASS]”).
-We denote the text attention matrix as $M\in\mathbb{R}^{1+N_{p}+N_{t}}$, where $N_{p}$ is the number of learnable prompts, $N_{t}$ is the length of input textual feature tokens.
+We denote the text attention matrix as $M\in\mathbb{R}^{(1+N_{p}+N_{t})*(1+N_{p}+N_{t})}$, where $N_{p}$ is the number of learnable prompts, $N_{t}$ is the length of input textual feature tokens.
 The additional 1 corresponds to the [EOS] token.
@@
 5.2 Results and Discussion
-This section compares GPR-NIAM-enhanced variants (GPR-NIAMP and GPR-NIAMT, derived by integrating GPR-NIAM into PromptFL and TCP, respectively) with nine baselines: Zero-Shot CLIP radford2021learning with hand-crafted text prompt template, e.g., “a photo of a [class]”, PromptFL guo2023PromptFL , RPO lee2023read , FedTPG qiu2024federated , TCP yao2024tcp , PromptFolio pan2024federated , DP-FPL tran2025privacy , FedDISC yang2024exploring , FedELMY wang2024one .
+This section compares GPR-NIAM-enhanced variants (GPR-NIAMP and GPR-NIAMT, derived by integrating GPR-NIAM into PromptFL and TCP, respectively) with nine baselines: Zero-Shot CLIP radford2021learning with hand-crafted text prompt template, e.g., “a photo of a [class]” li2024knn , PromptFL guo2023PromptFL , RPO lee2023read , FedTPG qiu2024federated , TCP yao2024tcp , PromptFolio pan2024federated , DP-FPL tran2025privacy , FedDISC yang2024exploring , FedELMY wang2024one .
 For fair comparison, all methods use the same backbone: a frozen CLIP with ViT-B16 radford2021learning .
@@
 Furthermore, in the Leave-One-Domain-Out Generalization task, similar conclusions are drawn, as GPR-NIAM outperforms the baseline across multiple domains.
+5.5 Performance Comparison across Different Local Training Epoch Settings
+Table 4: Performance comparison between GPR-NIAM and baselines on DTD and Flowers102 with local training epochs $\in\{1,5,20\}$ in the base-to-base/novel generalization setting.
+(a) DTD (Epochs=1) (b) DTD (Epochs=5) (c) DTD (Epochs=20)
+Method Base Novel HM Method Base Novel HM Method Base Novel HM
+CLIP 53.35 59.05 56.06 CLIP 53.35 59.05 56.06 CLIP 53.35 59.05 56.06
+PromptFL 50.34 51.69 51.01 PromptFL 55.55 56.76 56.15 PromptFL 63.31 51.08 56.54
+RPO 52.77 59.42 55.90 RPO 53.24 59.78 56.32 RPO 53.93 60.26 56.92
+FedTPG 40.27 36.47 38.28 FedTPG 49.88 47.22 48.51 FedTPG 56.25 56.28 56.26
+TCP 53.24 53.38 53.31 TCP 57.63 56.40 57.01 TCP 63.07 59.54 61.25
+PromptFolio 52.54 53.50 53.02 PromptFolio 58.68 56.28 57.45 PromptFolio 58.68 57.60 58.13
+DP-FPL 43.75 41.54 42.61 DP-FPL 59.25 50.12 54.30 DP-FPL 64.12 47.82 54.78
+GPR-NIAMP 65.97 51.20 57.65 GPR-NIAMP 68.51 52.65 59.54 GPR-NIAMP 70.13 53.26 60.54
+GPR-NIAMT 63.07 51.81 56.89 GPR-NIAMT 65.85 55.07 59.98 GPR-NIAMT 69.67 54.83 61.36
+(d) Flowers102 (Epochs=1) (e) Flowers102 (Epochs=5) (f) Flowers102 (Epochs=20)
+Method Base Novel HM Method Base Novel HM Method Base Novel HM
+CLIP 62.39 78.10 69.36 CLIP 62.39 78.10 69.36 CLIP 62.39 78.10 69.36
+PromptFL 70.84 74.10 72.43 PromptFL 78.42 76.00 77.19 PromptFL 80.46 73.68 76.92
+RPO 69.97 76.21 72.95 RPO 71.72 76.84 74.19 RPO 74.05 77.05 75.52
+FedTPG 68.51 77.05 72.53 FedTPG 73.76 77.05 75.37 FedTPG 74.34 77.05 75.67
+TCP 75.51 78.10 76.78 TCP 76.09 79.15 77.59 TCP 79.88 78.31 79.09
+PromptFolio 73.76 76.42 75.06 PromptFolio 79.59 75.57 77.53 PromptFolio 83.09 74.52 78.57
+DP-FPL 69.38 78.31 73.58 DP-FPL 73.76 79.57 76.55 DP-FPL 75.80 75.78 75.79
+GPR-NIAMP 82.50 75.57 78.89 GPR-NIAMP 84.83 78.31 81.44 GPR-NIAMP 85.13 77.47 81.12
+GPR-NIAMT 87.11 77.83 82.20 GPR-NIAMT 86.58 78.94 82.59 GPR-NIAMT 87.75 78.73 83.00
+To assess the robustness of the proposed GPR-NIAM under different local training efforts, we evaluate its performance with local epochs $\in{1,5,20}$, a data heterogeneity level of $\beta=0.5$, and 10 participating clients.
+As shown in Table 4, both GPR-NIAMP and GPR-NIAMT consistently outperform baselines across all datasets and settings.
+Notably, GPR-NIAMT achieves the highest harmonic mean (HM) in most cases, indicating better generalization across base and novel classes.
+On the Flowers102 dataset, it maintains strong performance with increasing local epochs, reflecting robustness against overfitting.
+Overall, both variants remain stable and effective despite changing local training dynamics.
 5.6 Effectiveness of Global Prompt Refinement
@@
 The communication overhead from prototype transmission may become a bottleneck in large-scale federated systems with limited bandwidth, which we plan to address in future work.
-And both the multi-label problem kou2025labeldistributionlearningbiased ; ijcai2024p478 and real-world medical imaging tasks feng2023unsupervised ; feng2025neighbor become considerably more challenging in federated settings.
+And both the multi-label kou2025labeldistributionlearningbiased ; ijcai2024p478 , multi-modal meng2025causal ; zhengwang_tnnls2025 ; wang2023multi ; cheng2023voice problems and real-world medical imaging tasks feng2023unsupervised ; feng2025neighbor ; dang2025deep become considerably more challenging in federated settings.
 In addition, several issues discussed in Section 5.8 remain open and deserve further investigation.
+Acknowledgments
+This work is supported in part by the Key Research and Development Program of Shandong Province (Grant No.
+2024TSGC0667); the Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RG101/24); the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).
 Appendix A Appendix
@@
 A.2.3 Performance Comparison across Different Local Training Epoch Settings
-Table 7: Performance comparison between GPR-NIAM and baselines on DTD and Flowers102 with local training epochs $\in\{1,5,20\}$ in the base-to-base/novel generalization setting.
-(a) DTD (Epochs=1) (b) DTD (Epochs=5) (c) DTD (Epochs=20)
-Method Base Novel HM Method Base Novel HM Method Base Novel HM
-CLIP 53.35 59.05 56.06 CLIP 53.35 59.05 56.06 CLIP 53.35 59.05 56.06
-PromptFL 50.34 51.69 51.01 PromptFL 55.55 56.76 56.15 PromptFL 63.31 51.08 56.54
-RPO 52.77 59.42 55.90 RPO 53.24 59.78 56.32 RPO 53.93 60.26 56.92
-FedTPG 40.27 36.47 38.28 FedTPG 49.88 47.22 48.51 FedTPG 56.25 56.28 56.26
-TCP 53.24 53.38 53.31 TCP 57.63 56.40 57.01 TCP 63.07 59.54 61.25
-PromptFolio 52.54 53.50 53.02 PromptFolio 58.68 56.28 57.45 PromptFolio 58.68 57.60 58.13
-DP-FPL 43.75 41.54 42.61 DP-FPL 59.25 50.12 54.30 DP-FPL 64.12 47.82 54.78
-GPR-NIAMP 65.97 51.20 57.65 GPR-NIAMP 68.51 52.65 59.54 GPR-NIAMP 70.13 53.26 60.54
-GPR-NIAMT 63.07 51.81 56.89 GPR-NIAMT 65.85 55.07 59.98 GPR-NIAMT 69.67 54.83 61.36
-(d) Flowers102 (Epochs=1) (e) Flowers102 (Epochs=5) (f) Flowers102 (Epochs=20)
-Method Base Novel HM Method Base Novel HM Method Base Novel HM
-CLIP 62.39 78.10 69.36 CLIP 62.39 78.10 69.36 CLIP 62.39 78.10 69.36
-PromptFL 70.84 74.10 72.43 PromptFL 78.42 76.00 77.19 PromptFL 80.46 73.68 76.92
-RPO 69.97 76.21 72.95 RPO 71.72 76.84 74.19 RPO 74.05 77.05 75.52
-FedTPG 68.51 77.05 72.53 FedTPG 73.76 77.05 75.37 FedTPG 74.34 77.05 75.67
-TCP 75.51 78.10 76.78 TCP 76.09 79.15 77.59 TCP 79.88 78.31 79.09
-PromptFolio 73.76 76.42 75.06 PromptFolio 79.59 75.57 77.53 PromptFolio 83.09 74.52 78.57
-DP-FPL 69.38 78.31 73.58 DP-FPL 73.76 79.57 76.55 DP-FPL 75.80 75.78 75.79
-GPR-NIAMP 82.50 75.57 78.89 GPR-NIAMP 84.83 78.31 81.44 GPR-NIAMP 85.13 77.47 81.12
-GPR-NIAMT 87.11 77.83 82.20 GPR-NIAMT 86.58 78.94 82.59 GPR-NIAMT 87.75 78.73 83.00
 Table 8: Performance comparison between GPR-NIAM and baselines on StanfordCars and UCF101 with local training epochs $\in\{1,5,20\}$ in the base-to-base/novel generalization setting.
@@
 Specifically, we consider three settings with local training epochs $\in\{1,5,20\}$, under a data heterogeneity level ($\beta=0.5$) and with 10 clients participating in training.
-As shown in Table 7 and Table 8, both GPR-NIAMP{P} and GPR-NIAMT{T} consistently outperform the baselines across all datasets and training epochs.
+As shown in Table 4 and Table 8, both GPR-NIAMP and GPR-NIAMT consistently outperform the baselines across all datasets and training epochs.
 Notably, GPR-NIAMT{T} delivers the best overall performance in terms of harmonic mean (HM) in most settings, which demonstrates superior generalization across both base and novel classes.
@@
 Across datasets, both variants of GPR-NIAM remain stable and effective even as the local training dynamics change, outperforming methods that are more sensitive to overfitting or class imbalance under prolonged local updates.
-A.2.4 Evaluating Method Scalability with Varying Client Numbers
-Table 9: Performance comparison between GPR-NIAM and baselines on DTD and Flowers102 datasets under varying number of clients $K\in\{20,30\}$ in the base-to-base/novel generalization setting.
+A.2.4 Sensitivity Analysis Regarding the Number of Prompt Tokens.
+Table 9: Comparison of different methods under two settings (5×512 and 20×512) on Caltech101, DTD, and UCF101 datasets. Bold numbers indicate the best performance.
+Caltech101 DTD UCF101
+Base Novel HM Base Novel HM Base Novel HM
+5×512
+PromptFL 88.41 93.97 91.11 59.25 52.77 55.83 76.78 69.91 73.18
+FedTPG 86.65 93.37 89.88 60.18 56.15 58.10 74.30 70.63 72.41
+PromptFolio 89.47 93.46 91.42 61.02 56.28 57.62 74.40 70.65 72.48
+GRP-NIAM_P 89.90 93.37 91.60 60.96 55.95 59.12 77.98 70.39 73.99
+20×512
+PromptFL 88.91 94.31 91.53 62.50 52.53 57.08 76.83 69.12 72.77
+FedTPG 88.84 92.86 90.81 58.10 54.83 56.41 76.30 73.41 74.41
+PromptFolio 89.97 94.22 92.05 56.13 53.98 55.03 75.72 74.41 75.06
+GRP-NIAM_P 90.02 94.28 92.10 61.43 56.23 58.71 77.23 74.58 75.88
+We have conducted additional sensitivity experiments regarding the number of prompt tokens.
+Specifically, we adjusted the prompt length to 5×512 and 20×512.
+The results show that GPR-NIAM remains robust across different prompt lengths, while consistently maintaining an advantage over all baselines in all case.
+A.2.5 Evaluating Method Scalability with Varying Client Numbers
+Table 10: Performance comparison between GPR-NIAM and baselines on DTD and Flowers102 datasets under varying number of clients $K\in\{20,30\}$ in the base-to-base/novel generalization.
 (a) DTD (K=20) (b) DTD (K=30)
@@
 GPR-NIAMT 83.96 77.05 80.36 GPR-NIAMT 82.50 78.94 80.68
-Table 10: Performance comparison between GPR-NIAM and baselines on StanfordCars and UCF101 datasets under varying number of clients $K\in\{20,30\}$ in the base-to-base/novel generalization setting.
+Table 11: Performance comparison between GPR-NIAM and baselines on StanfordCars and UCF101 datasets under varying number of clients $K\in\{20,30\}$ in the base-to-base/novel generalization.
 (g) StanfordCars (K=20) (h) StanfordCars (K=30)
@@
 These results show that GPR-NIAM maintains high performance and stability even as the number of clients increases, demonstrating its scalability and effectiveness in more challenging cases.
-A.2.5 Evaluation of Module Contributions
+A.2.6 Evaluation of Module Contributions
 Table 12: Ablation study on the effectiveness of NIAM and CSCR modules under the base-to-base/novel generalization. Results are reported on four datasets with PromptFL as the baseline.
@@
 Together, they deliver the best overall performance in terms of harmonic mean (HM), highlighting their complementary effects.
+A.2.7 Sensitive Analysis of Textual Prompt Template
+Table 13: Prompt template comparison on Caltech101, DTD, and UCF101. Best numbers in bold.
+Template Caltech101 DTD UCF101
+Base Novel HM Base Novel HM Base Novel HM
+A photo of a [CLASS] 91.59 94.22 92.89 68.40 53.50 60.04 78.57 70.28 74.19
+A picture of a [CLASS] 91.43 94.34 92.86 67.37 52.83 59.22 78.61 69.54 73.79
+An image showing [CLASS] 91.52 93.88 92.68 68.57 54.12 60.49 78.77 70.14 74.20
+The model’s performance is not highly sensitive to the choice of textual prompt templates.
+Following prior works such as CLIP and PromptFL, we use “a photo of a [CLASS]” as the default template.
+Since our method mainly optimizes the learnable prompt tokens ($T_{P}$) while keeping the text tokens ($T_{T}$) fixed, the initial template primarily serves as a starting point for semantic alignment and has limited impact on the final performance.
+We also conducted a small-scale experiment comparing different templates (e.g., “a picture of a [CLASS]”, “an image showing [CLASS]”), and the variation in Top-1 accuracy was less than 1%, indicating that the GPR-NIAM framework is robust to the choice of templates.
 NeurIPS Paper Checklist